Justin is 4 years older than Ethan. In 9 years the sum of their ages will be 28. Find the age of each boy now.
Found 2 solutions by PRMath, solver91311:
Answer by PRMath(133)

Justin is 4 years older than Ethan. In 9 years the sum of their ages will be 28. Find the age of each boy now.
Justin is 4 years older than Ethan. Justin will ALWAYS be 4 years older than Ethan.
So let's say Ethan is "E"
Justin is "J".
Some day (we don't know when) but some day, when you add their ages together, their ages will add up to be 28.
So, we can say: E + J = 28
But.... Justin is 4 years older than Ethan: J = 4 + E

Therefore, our equation of E + J = 28 should REALLY be:
J = 4 + E
E + 4 + E = 28

Do you see how we replaced "J" with "4 + E"?
Ok, so now we have:

E + 4 + E = 28 (Let's combine our like terms)
2E + 4 = 28 (1E + 1E = 2E)
2E = 28 - 4 (Let's subtract 4 from both sides)
2E = 24 (Now let's divide both sides by 2)
E = 12
SO if Ethan is 12 and Justin is 4 years older, Justin is 16.

ETHAN = 12
JUSTIN = 12 + 4 which is 16
If you combine (add) their ages: 12 + 16 = 28
BUT they are at that age in NINE years! So what are they now? Just subtract 9 from both of their ages:
Ethan: 12 - 9 = 3
Justin: 16 - 9 = 7
At 7, Justin is 4 years older than Ethan, who is three.
SOooooo right now they are 3 and 7.

Answer by solver91311(24713)



Let represent Ethan's age now. Then is Justin's age now. Nine years from now, Ethan will be and Justin will be



Solve for

John